At a prayer meeting the disciples sing JAI RAM JAI RAM the sound amplified by a loudspeaker comes back after reflection from a building at a distance of 80 m from the meeting. What max time  intervals can be kept between one JAI RAM and the next JAI RAM so that the echo does not disterb a listener sitting in the meeting. Speed of sound in air is 320m/s.

Grade:12

2 Answers

Abhishek Kumar
8 years ago
Time taken to go to the wall = 80/320 s = Time taken to come back
So, Time difference = 2 x 80/320 = 0.5 Sec
Leonard
8 years ago
distance between speaker and building=80m
total distance=80x2=160m
given v=320
t=s/v
t=160/320=0.5 sec